# Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs

### v1

@@ 3.1.2 Ranked Evaluation
$Rank(m)=\sum_{c\in C}S(m_{k},c)$ (1)
Tables 3 and 4 present the results of this ranked evaluation. The ranked evaluation aligns with the Best-Performer Evaluation in identifying Qwen 2.5 14B as the most dominant model. However, it also offers insights into the relative performance of other models, highlighting distinctions among high performers.
@@ 3.2 Parameter Performance Analysis
In this section, we examine how different parameter configurations drive model performance. Identifying the most impactful parameters was not a straightforward task due to the complexity of the search space and the variety of configurations tested. To address this, we conducted a Quantile-Based Analysis, focusing on the top $5$% of configurations based on F1-scores. Examining the distribution of parameter settings within this high-performing subset, allows us to uncover trends and patterns that could guide the selection of optimal configurations.
For this analysis, we focused exclusively on the SciCite dataset, and the insights gained from it were then used to limit the range of experiments conducted on the ACL-ARC dataset. The findings from this analysis can be seen in Table 5 and are summarized below.
Rank Model Score
1 Qwen 2.5 – 14B 26.000
2 Gemma 2 – 27B 9.747
3 Gemma 2 – 9B 8.498
4 Qwen 2 – 7B 7.554
5 LLaMA 3.1 – 8B 6.437
6 Mistral Nemo – 12B 6.000
7 Phi 3.5 Mini – 3.8B 5.268
8 Phi 3 Medium - 14B 3.399
9 LLaMA 3.2 - 3B 3.281
10 LLaMA 3.2 - 1B 2.718
11 Gemma 2 - 2B 2.506
12 LLaMA 3 - 8B 2.396
Table 4: Model ranking based on RRF on ACL-ARC.
We discern that few-shot prompting was the most common prompting method among the top-performing configurations, accounting for nearly 45% of the results. Many-shot prompting followed closely with 34.83%, while one-shot and zero-shot prompting were significantly less represented, with only 16.85% and 3.37%, respectively. This suggests that providing several examples during prompting generally leads to better performance, with the optimal achieved at five examples (few-shot), followed by a slight decrease at larger numbers of examples.
In contrast, the temperature parameter exhibited a relatively even distribution among the top-performing configurations, with $0.0$ and $0.2$ being the most frequent settings (30.34% each). Higher temperatures, such as $1.0$ and $0.5$, were less common, suggesting that lower temperatures contribute to more consistent and deterministic outputs, which align better with the task requirements.
Similarly, the system prompt parameter was fairly evenly distributed among the top configurations, with SP3 being slightly more frequent (35.96%), followed by SP2 (32.58%), and SP1 (31.46%). While all system prompts contributed to high performance, these results suggest that SP3 has offered a slight advantage in effectiveness.
Parameter Setting Count Percent
Prompting Method Few-shot 40 44.94%
Many-shot 31 34.83%
One-shot 15 16.85%
Zero-shot 3 3.37%
Temperature 0.0 27 30.34%
0.2 27 30.34%
1.0 18 20.22%
0.5 17 19.10%
System Prompt SP3 32 35.96%
SP2 29 32.58%
SP1 28 31.46%
Query Template Multiple-Choice 83 93.26%
Simple 6 6.74%
Examples Method Roles 44 51.16%
Inline 42 48.84%
Table 5: Parameter Performance Analysis of Top 5% Configurations on SciCite (sorted by percentage).
The query template parameter showed a strong preference for multiple-choice templates, which appeared in 93.26% of the top-performing configurations. This overwhelming dominance suggests that multiple-choice queries are significantly more effective for the evaluated task, likely due to their structured.
The examples method parameter showed a near-equal split between roles (51.16%) and inline (48.84%), indicating that both approaches are effective, with no clear advantage for either method among the top-performing configurations.
@@
These findings offer valuable insights into the parameter settings that consistently enhance performance. To verify this analysis, we also performed a Chi-Square statistical test, outlined in Appendix B.
@@ 4 Fine-tuning
Parameter Value
Learning Rate 5e-5
Epochs 10
Batch Size 16
Cutoff Length 512
Optimizer AdamW
Compute Type fp16
Warmup Steps 500
DeepSpeed Offload Enabled
DeepSpeed Stage 3
LoRA Rank 8
LoRA Alpha 16
LoRA Dropout 0.1
@@
4.1 Training Configuration
For this experiment, we used the SciCite and ACL-ARC datasets, which were converted into the Alpaca format (Taori et al., 2023). This format includes a system prompt, an instruction, the citing sentence, and the true label. The use of Supervised Fine-Tuning (SFT) was motivated by its ability to adapt pre-trained models to specific tasks using minimal labeled data, making it an effective approach for citation intent classification.
To fine-tune the Qwen 2.5 14B Instruct model, we used LLaMA-Factory (Zheng et al., 2024) on an AWS g6e.12xlarge EC2 instance equipped with 4 NVIDIA L40S GPUs, providing a combined GPU memory of 192GB. The training process was conducted using fp16 mixed precision to optimize memory usage and speed. To prevent memory issues during fine-tuning, we employed DeepSpeed ZeRO Stage 3 Offload (Ren et al., 2021), which enabled efficient memory management by offloading optimizer states and gradients to the CPU. The training parameters included a learning rate of 5e-5, a batch size of 16, and 10 epochs. The model was optimized using AdamW, with a cutoff length of 512 tokens for input sequences.
To optimize the fine-tuning process, we employed Low-Rank Adaptation (LoRA), a widely used parameter-efficient fine-tuning (PEFT) method (Hu et al., 2021; Han et al., 2024). LoRA enables efficient adaptation of large language models by freezing the pre-trained model weights and introducing trainable low-rank matrices into the Transformer layers, significantly reducing the number of trainable parameters while maintaining performance. For this experiment, we configured LoRA with a rank of 8, alpha of 16, and a dropout of 0.1, which allowed us to fine-tune the model effectively without exceeding memory constraints. The outlined training parameters are also summarized in Table 6.
@@ 4.2 Results
After fine-tuning, we evaluated the new model using the same prompting methods as in Section 2.3.1 (i.e., zero-shot, one-shot, few-shot, and many-shot) to draw comparisons with the baseline Qwen 2.5 14B Instruct model. Although in the previous evaluation we only used 8-bit quantized models (Q8) due to the extensiveness of our experiment space, in this experiment, we also evaluate the 16-bit floating-point (FP16) version of the model to examine the impact of bit precision on performance.
@@
Instruct Q8 78.33 61.04
Instruct FP16 80.41 61.73
Fine-tuned Q8 86.47 68.48
Fine-tuned FP16 86.84 67.73
Table 7: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants on the SciCite and ACL-ARC datasets.
The results in Table 7 show that fine-tuning improved performance on both SciCite and ACL-ARC. On SciCite, the fine-tuned FP16 model achieved an F1-score of 86.84, an 8% improvement over the instruction-tuned baseline, while the fine-tuned Q8 model achieved a comparable score of 86.47. On ACL-ARC, the fine-tuned Q8 model achieved the highest F1-score of 68.48, outperforming the instruction-tuned baseline by nearly 11%.
Precision settings showed minor differences, with FP16 models generally outperforming Q8 on SciCite, while Q8 slightly outperformed FP16 on ACL-ARC. This suggests that while FP16 precision is generally advantageous, Q8 remains competitive and may offer robustness for tasks with more granular classification schemes, such as ACL-ARC’s 6-class setup.
Table 8 highlights the impact of fine-tuning across prompting methods. On SciCite, fine-tuning led to consistent improvements, particularly in zero-shot and one-shot settings, where the fine-tuned Q8 model improved by nearly 10% over the instruction-tuned baseline. In few-shot and many-shot setups, performance approached saturation, with the FP16 model achieving the highest F1-scores of 86.84 and 85.79, respectively. On ACL-ARC, fine-tuned models showed similar trends, with the Q8 model achieving the highest F1-scores in one-shot (68.48) and few-shot (64.35) setups, outperforming FP16 in most cases. These results demonstrate that fine-tuning enhances generalization in low-context scenarios while effectively leveraging additional context during inference.
Dataset Model Zero- One- Few- Many-
Shot Shot Shot Shot
SciCite Instruct Q8 75.26 76.29 78.33 78.32
@@
ACL-ARC Instruct Q8 50.02 60.02 61.04 59.75
Instruct FP16 50.17 61.73 59.04 54.46
Fine-tuned Q8 59.40 68.48 64.35 65.43
Fine-tuned FP16 60.00 67.73 64.34 64.62
Table 8: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants on the SciCite and ACL-ARC datasets divided by prompting method.
@@ 4.3 Discussion
We consider a key advantage of LLMs to be their ease of use and deployment. Tools such as LM Studio and Ollama [^5] https://ollama.com/ allow models like those outlined in this paper to be deployed locally with zero technical expertise, making them accessible to users without a computer science background. In addition to their accessibility, LLMs offer significant adaptability. Unlike traditional methods, which often require complex pretraining or domain-specific tuning, LLMs can be fine-tuned for a wide range of scientometric tasks, such as citation recommendation, paper summarization, or trend analysis, without the need for bespoke architectures. Furthermore, LLMs can scale effectively to new scientific domains or datasets, requiring only small amounts of task-specific data to adapt to underexplored scientific fields. For instance, our fine-tuned models were trained only on the several thousand citing sentences provided by our task-specific datasets, whereas SciBERT has been trained on millions of scientific papers.
@@ 5 Related Work
The theoretical beginnings of citation analysis can be traced back to foundational works such as Garfield et al.’s (1965) identification of reasons for citation Moravcsik and Murugesan’s (1975) studies on citation function. Early annotation schemes, such as those by Spiegel-Rosing (1977), were later adapted by Teufel et al. (2006) for supervised machine learning approaches to citation classification.
Jurgens et al. (2018) introduced the ACL-ARC dataset, which contains nearly $2,000$ citations from papers in the NLP field, annotated for their function with a classification scheme of six classes. Pride and Knoth (2020) extended this classification scheme by refining the comparison class to capture similarities, differences, and disagreement. Around the same time, Cohan et al. (2019) proposed a multitask model incorporating structural information from scientific papers. They also introduced the SciCite dataset, which is significantly larger and spans multiple scientific domains, with three intent classes.
Beltagy et al. (2019) introduced SciBERT, a BERT-based (Devlin et al., 2019) pre-trained language model (PLM) for scientific text, which has since become the backbone of many citation intent classification methods. SciBERT has been widely adopted due to its ability to generalize across scientific domains. For example, Mercier et al. (2021) introduced ImpactCite, an XLNet-based method for citation impact analysis, which was later used by Paolini et al. (2024) to achieve state-of-the-art results on the SciCite dataset. Paolini et al. demonstrated the effectiveness of ensemble classifiers combining fine-tuned SciBERT and XLNet models. On the ACL-ARC dataset, Hu et al. (2022) achieved state-of-the-art results by studying domain adaptation of language models using a variational masked autoencoder.
@@
Feature-rich RF (Jurgens et al., 2018) 53.00
Structural Scaffolds (Cohan et al., 2019) 84.00 67.90
SciBERT (Beltagy et al., 2019) 85.22
ImpactCite (Mercier et al., 2021) 88.93
VarMAE Hu et al. (2022) 86.32 76.50
CitePrompt (Lahiri et al., 2023) 86.33 68.39
EnsIntWS (Paolini et al., 2024) 89.46
EnsIntWoS (Paolini et al., 2024) 88.48
MTL Finetuning (Search) (Shui et al., 2024) 85.25 64.56
MTL Finetuning (TRL) (Shui et al., 2024) 85.35 75.57
Recent research has continued to explore PLM-based methods for citation intent classification. Lahiri et al. (2023) used a prompt-based learning approach on SciBERT to identify citation intent, while Shui et al. (2024) propose a multi-task learning framework that jointly fine-tunes SciBERT on a dataset of primary interest together with multiple auxiliary datasets to take advantage of additional supervision signals. Kunnath et al. (2023) explored various prompting and tuning strategies on SciBERT, including fixed and dynamic context prompts, and found that parameter updating with prompts improved performance. They also briefly experiment on LLMs by evaluating the zero-shot performance of GPT-3.5, which performed well on their recently introduced ACT2 dataset but poorly on the ACL-ARC dataset. However, GPT-3.5 was not evaluated on SciCite. In contrast, our work is the first to focus entirely on evaluating and fine-tuning numerous open-weight LLMs without using models pre-trained specifically for the scientific domain.

### v2

@@ 3 Gemma 2 – 9B 11 5 5 1 0
$RankingScore(m)=\sum_{c\in C}S(m_{k},c)$ (1)
Table 3 presents the results of this ranked evaluation. The ranked evaluation aligns with the Best-Performer Evaluation in identifying Qwen 2.5 14B as the most dominant model. This view also allows us to gain insights into the relative performance of other the models, highlighting distinctions among high performers, as well as the performance of lower-ranked models.
Table 4: Parameter Performance Analysis of Top 5% Configurations.
Parameter Setting SciCite ACL-ARC
Count Percent Count Percent
In-context Learning Method Few-shot 47 46.53% 55 61.11%
Many-shot 38 35.64% 21 23.33%
One-shot 14 13.86% 14 15.55%
Zero-shot 4 3.96% - -
Temperature 0.0 31 30.69% 23 25.55%
0.2 30 29.70% 22 24.44%
0.5 20 19.80% 22 24.44%
1.0 20 19.80% 23 25.55%
System Prompt SP3 45 44.55% 38 42.22%
SP2 28 27.72% 24 26.66%
SP1 28 27.72% 28 31.11%
Query Template Multiple-Choice 80 79.21% 58 64.44%
Simple 21 20.79% 32 35.55%
Example Presentation Conversational 53 54.64% 59 65.55%
Inline 44 45.36% 31 34.44%
@@ 3.2 Parameter Performance Analysis
In this section, we examine how different parameter configurations drive model performance. Identifying the most impactful parameters was challenging due to the extensive search space ($168$ configurations per dataset, totaling $3,841$ experiments for all models). To address this, we conducted a Quantile-Based Analysis focusing on the top $5$% of configurations based on F1-scores. Examining parameter distributions within this high-performing subset (Table 4) reveals trends for optimal configurations across both datasets.
For the In-context Learning Method, Few-shot prompting was most prevalent in top SciCite configurations ($46.53$%), with Many-shot also prominent ($35.64$%). This Few-shot preference was stronger on ACL-ARC ($61.11$%), where Many-shot ($23.33$%) and One-shot ($15.55$%) were less frequent, and Zero-shot was absent from the top 5%. These results suggest moderate example counts (Few-shot) are broadly effective, though Many-shot also performs well on SciCite.
The optimal Temperature settings showed divergence between the datasets. SciCite favored lower values, with T=0.0 ($30.69$%) and T=0.2 ($29.70$%) being most frequent, indicating a preference for deterministic outputs, while higher temperatures (T=0.5, T=1.0, both $19.80$%) were less common. In contrast, ACL-ARC displayed a more balanced distribution, where T=0.0 and T=1.0 (both $25.55$%) were slightly ahead of T=0.2 and T=0.5 (both $24.44$%), suggesting that varied levels of randomness can yield top results.
A shared preference for System Prompt SP3 emerged across both datasets, found in $44.55$% of top SciCite and $42.22$% of top ACL-ARC configurations. Secondary preferences, however, varied. SciCite showed equal representation for SP1 and SP2 (both $27.72$%), whereas ACL-ARC favored SP1 ($31.11$%) over SP2 ($26.66$%). This indicates an advantage for the structured SP3, though simpler prompts also achieved high performance.
The Query Template parameter revealed a decisive trend, as the Multiple-Choice template was overwhelmingly favored on both SciCite ($79.21$%) and ACL-ARC ($64.44$%) when compared to the Simple template (SciCite: $20.79$%; ACL-ARC: $35.55$%). This highlights the significant effectiveness of explicit multiple-choice options.
Finally, considering the Example Presentation Format, the Conversational style was more frequent in top configurations for SciCite ($54.64$%) and especially ACL-ARC ($65.55$%), over the Inline format (SciCite: $45.36$%; ACL-ARC: $34.44$%). This suggests structuring examples conversationally generally yields better results.
@@
These findings offer valuable insights into the parameter settings that consistently enhance performance across the two datasets.
Table 5: Chi-Square Test of Independence results for parameter settings and F1-Scores. (Significance levels: ***p < 0.001, **p < 0.01, *p < 0.05)
Parameter SciCite ACL-ARC
Chi-Square ($\chi^{2}$) p-value Chi-Square ($\chi^{2}$) p-value
In-context Learning Method 29.857 1.48e-06 *** 49.025 1.29e-10 ***
Temperature 4.617 0.202 0.047 0.9973
System Prompt 6.024 0.0491 * 3.647 0.1615
Query Template 35.063 3.19e-09 *** 7.305 0.0068 **
Example Presentation 0.699 0.4031 8.604 0.0033 **
3.2.1 Chi-Square Statistical Test.
To complement our parameter performance analysis, we also conducted a Chi-Square Test of Independence [12] to evaluate the relationship between parameter settings and F1-scores. This approach allowed us to determine whether specific parameter settings were significantly associated with performance outcomes and validate the results of our primary analysis. The results of the test (summarized in Table 5) largely reinforces the findings from our quantile-based analysis regarding the influence of different parameter settings.
The In-context Learning Method showed a highly significant association with F1-scores for both SciCite ($p<0.001$) and ACL-ARC ($p<0.001$). This statistically underscores the importance of the number of examples provided, aligning with our earlier observation that Few-shot and Many-shot configurations were prevalent among the top performers.
Similarly, the Query Template was found to be highly significant for SciCite ($p<0.001$) and also significant for ACL-ARC ($p<0.01$). This supports the strong dominance of the Multiple-Choice template in the top 5% configurations, highlighting its critical role in guiding model output effectively.
Other parameters presented more nuanced relationships. The System Prompt was statistically significant for SciCite ($p<0.05$), corroborating the advantage seen for SP3 in its quantile analysis. However, for ACL-ARC, while SP3 was also frequent in top configurations, the choice of system prompt did not emerge as a statistically significant factor overall.
Conversely, the Example Presentation Format was not statistically significant for SciCite, which aligns with the relatively balanced distribution (Conversational at 54.64%, Inline at 45.36%) observed in its top configurations. For ACL-ARC, however, this parameter was significant ($p<0.01$), providing statistical backing for the clearer preference (65.55%) for the Conversational method seen in its quantile analysis.
Finally, the Temperature parameter did not show a statistically significant relationship with F1-scores on either dataset. This finding is consistent with the quantile analysis where, for SciCite, multiple temperature settings were present in top configurations (though lower ones were more frequent), and for ACL-ARC, the distribution was particularly balanced, suggesting that temperature, within the tested ranges, was not as decisive a factor as other parameters.
@@ 4 Fine-tuning
4.1 Training Configuration
For this experiment, we used the SciCite and ACL-ARC datasets, which were converted into the Alpaca format [30]. This format includes a system prompt, an instruction, the citing sentence, and the true label. The use of Supervised Fine-Tuning (SFT) was motivated by its ability to adapt pre-trained models to specific tasks using minimal labeled data, making it an effective approach for citation intent classification.
To fine-tune the Qwen 2.5 14B Instruct model, we used LLaMA-Factory [33] on an AWS g6e.12xlarge EC2 instance equipped with 4 NVIDIA L40S GPUs, providing a combined GPU memory of 192GB. The training process was conducted using fp16 mixed precision to optimize memory usage and speed. To prevent memory issues during fine-tuning, we employed DeepSpeed ZeRO Stage 3 Offload [27], which enabled efficient memory management by offloading optimizer states and gradients to the CPU. The training parameters included a learning rate of 5e-5, a batch size of 16, and 10 epochs. The model was optimized using AdamW, with a cutoff length of 512 tokens for input sequences.
To optimize the fine-tuning process, we employed Low-Rank Adaptation (LoRA), a widely used parameter-efficient fine-tuning (PEFT) method [14, 13]. LoRA enables efficient adaptation of large language models by freezing the pre-trained model weights and introducing trainable low-rank matrices into the Transformer layers, significantly reducing the number of trainable parameters while maintaining performance. For this experiment, we configured LoRA with a rank of 8, alpha of 16, and a dropout of 0.1, which allowed us to fine-tune the model effectively without exceeding memory constraints. The outlined training parameters are also summarized in Table 6.
@@ 4.1 Training Configuration
Parameter Value Explanation
Learning Rate 5e-5 Controls the step size for updating model weights during training.
Epochs 10 Number of complete passes through the entire training dataset.
Batch Size 16 Number of training examples processed in one iteration.
Cutoff Length 512 Maximum sequence length (in tokens) for model inputs.
Optimizer AdamW Algorithm used to adjust model weights to minimize loss.
Compute Type fp16 Numerical precision used for computations.
Warmup Steps 500 Number of initial steps where the learning rate gradually increases.
DeepSpeed Offload Enabled Moves optimizer states/gradients to CPU RAM to save GPU memory.
DeepSpeed Stage 3 Level of DeepSpeed ZeRO optimization.
LoRA Rank 8 Dimension of the trainable low-rank matrices added by LoRA.
LoRA Alpha 16 Scaling factor for the LoRA updates relative to the rank.
LoRA Dropout 0.1 Dropout probability applied to the LoRA layers for regularization.
@@ 4.2 Results
After fine-tuning, we evaluated the new model using the same prompting methods as in Section 2.3.1 (i.e., zero-shot, one-shot, few-shot, and many-shot) to draw comparisons with the instruction-tuned Qwen 2.5 14B baseline. While our initial extensive experiments with in-context learning focused on 8-bit quantized models (Q8), for this fine-tuning evaluation, we also assessed the 16-bit floating-point (FP16) version of the model to examine the impact of numerical precision on performance.
The overall results, presented in Table 7, clearly show that supervised fine-tuning significantly improved performance on both the SciCite and ACL-ARC datasets. On SciCite, the fine-tuned FP16 model achieved a peak F1-score of $86.84$. Compared to the instruction-tuned FP16 baseline (F1-score of 80.41), this represents an $8.0$% relative improvement. On ACL-ARC, the fine-tuned Q8 model obtained the highest F1-score of $68.48$. This constitutes a relative improvement of nearly 4.3% over the instruction-tuned FP16 baseline (F1-score of $65.64$). Overall, the highest F1-score achieved after fine-tuning was $86.84$ on SciCite and $68.48$ on ACL-ARC.
Examining the impact of numerical precision, the fine-tuned FP16 and Q8 models showed minor performance differences. FP16 models generally outperformed Q8 on SciCite, while the fine-tuned Q8 model slightly outperformed its FP16 counterpart on ACL-ARC. This suggests that while FP16 precision can offer a slight advantage, Q8 remains highly competitive and may offer robustness, particularly for tasks with more granular classification schemes like ACL-ARC’s 6-class setup.
Table 8 highlights the impact of fine-tuning across the different prompting methods. On SciCite, fine-tuning led to consistent F1-score improvements across all prompting scenarios. These gains were particularly substantial in zero-shot and one-shot settings; for instance, the fine-tuned Q8 model improved upon the instruction-tuned FP16 baseline by over $12.5$% in the zero-shot setting ($84.84$ vs $75.38$) and by nearly $10.7$% in the one-shot setting ($85.46$ vs $77.22$). In the few-shot and many-shot setups for SciCite, performance with the fine-tuned FP16 model approached saturation, achieving the highest F1-scores of $86.84$ and $85.79$, respectively.
Similar trends were observed on the ACL-ARC dataset. The fine-tuned models generally outperformed their instruction-tuned counterparts across prompting methods, with the fine-tuned Q8 model achieving the highest F1-scores in one-shot ($68.48$) and few-shot ($67.58$) setups. In contrast, the fine-tuned FP16 model yielded stronger results in the zero-shot ($60.29$) and many-shot ($67.62$) scenarios.
These results demonstrate that supervised fine-tuning not only boosts overall performance but also particularly enhances generalization in low-context (zero-shot and one-shot) scenarios, while still effectively leveraging the additional context provided in few-shot and many-shot settings.
Table 7: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants.
@@
Qwen 2.5 – 14B Instruct Q8 78.33 63.68
Qwen 2.5 – 14B Instruct FP16 80.41 65.64
Qwen 2.5 – 14B Fine-tuned Q8 86.47 68.48
Qwen 2.5 – 14B Fine-tuned FP16 86.84 67.73
Table 8: F1-Score Performance of Qwen 2.5 – 14B Instruct and Fine-tuned variants, divided by prompting method.
Dataset Model Zero-Shot One-Shot Few-Shot Many-Shot
SciCite Instruct Q8 75.74 76.32 78.33 78.23
@@
ACL-ARC Instruct Q8 52.29 60.02 62.86 63.68
Instruct FP16 52.47 61.73 65.64 64.39
Fine-tuned Q8 59.92 68.48 67.58 67.19
Fine-tuned FP16 60.29 67.73 66.94 67.62
@@ 4.3 Discussion
We consider a key advantage of LLMs to be their ease of use and deployment. Tools such as LM Studio and Ollama [^8] https://ollama.com/ allow models like those outlined in this paper to be deployed locally with zero technical expertise, making them accessible to users without a computer science background. In addition to their accessibility, LLMs offer significant adaptability. Unlike traditional methods, which often require complex pretraining or domain-specific tuning, LLMs can be fine-tuned for a wide range of scientometric tasks, such as citation recommendation, paper summarization, or trend analysis, without the need for bespoke architectures. Furthermore, LLMs can scale effectively to new scientific domains or datasets, requiring only small amounts of task-specific data to adapt to underexplored scientific fields. , our fine-tuned models required only the several thousand citing sentences provided by our task-specific datasets for adaptation. This contrasts sharply with models like SciBERT, whose effectiveness stems from deliberate pre-training exclusively on a large corpus of scientific papers (millions of articles), optimizing it for scientific language understanding. While the general-purpose LLMs used in our study were likely exposed to scientific text within their vast, diverse pre-training data drawn from the web, this exposure is incidental rather than targeted. They were not specifically pre-trained or architected with the primary goal of processing scientific literature, unlike SciBERT. Our results demonstrate that even without such specialized scientific pre-training, general-purpose LLMs can achieve competitive performance through efficient fine-tuning on minimal task-specific data.
@@ 5 Related Work
The theoretical beginnings of citation analysis can be traced back to foundational works such as [9] identification of reasons for citation [23] studies on citation function. Early annotation schemes, such as those by [29], were later adapted by [32] for supervised machine learning approaches to citation classification.
[17] introduced the ACL-ARC dataset, which contains nearly $2,000$ citations from papers in the NLP field, annotated for their function with a classification scheme of six classes. [25] extended this classification scheme by refining the comparison class to capture similarities, differences, and disagreement. Around the same time, [5] proposed a multitask model incorporating structural information from scientific papers. They also introduced the SciCite dataset, which is significantly larger and spans multiple scientific domains, with three intent classes.
[3] introduced SciBERT, a BERT-based [7] encoder language model pre-trained specifically on scientific text, which has since become the backbone of many citation intent classification methods. SciBERT has been widely adopted due to its ability to generalize across scientific domains. For example, [22] introduced ImpactCite, an XLNet-based method for citation impact analysis, which was later used by [24] to achieve state-of-the-art results on the SciCite dataset. Paolini et al. demonstrated the effectiveness of ensemble classifiers combining fine-tuned SciBERT and XLNet models.
Recent research has continued to explore PLM-based methods for citation intent classification. [20] used a prompt-based learning approach on SciBERT to identify citation intent, while [28] achieved state-of-the-art performance on the ACL-ARC dataset by proposing a multi-task learning framework that jointly fine-tunes SciBERT on a dataset of primary interest together with multiple auxiliary datasets to take advantage of additional supervision signals. [19] explored various prompting and tuning strategies on SciBERT, including fixed and dynamic context prompts, and found that parameter updating with prompts improved performance. They also briefly experiment on LLMs by evaluating the zero-shot performance of GPT-3.5, which performed well on their recently introduced ACT2 dataset but poorly on the ACL-ARC dataset. However, GPT-3.5 was not evaluated on SciCite. In contrast, our work is the first to focus entirely on evaluating and fine-tuning numerous open-weight, general-purpose LLMs without using models pre-trained explicitly and exclusively for the scientific domain.
@@
Feature-rich Random Forest [17] 53.00
Structural Scaffolds [5] 84.00 67.90
SciBERT [3] 85.22
ImpactCite [22] 88.93
CitePrompt [20] 86.33 68.39
EnsIntWS [24] 89.46
EnsIntWoS [24] 88.48
MTL Finetuning (Search) [28] 85.25 64.56
MTL Finetuning (TRL) [28] 85.35 75.57